OpenStudy (samara8954):

In January, the depth of a lake was 833 feet. In August, the depth of the lake was 499.8 feet. What is the percentage decrease of the depth of the lake from January to August? 40% 48% 60% 67%

OpenStudy (samara8954):

@ms.hi

OpenStudy (samara8954):

@AlexandervonHumboldt2

OpenStudy (samara8954):

@dtuniyants

OpenStudy (samara8954):

are you sure?

OpenStudy (samara8954):

okay

OpenStudy (anonymous):

use formula: new/old*100%

OpenStudy (samara8954):

what is that

OpenStudy (anonymous):

499.8/833*100%

OpenStudy (anonymous):

you will find what persantage is 499.8 out of 833

OpenStudy (samara8954):

so the answer is 60

OpenStudy (anonymous):

let me calculate.

OpenStudy (samara8954):

am I correct 60

OpenStudy (anonymous):

499.8 is 60% from 833. so the decrease is 40%

OpenStudy (anonymous):

i was right in the beggining

OpenStudy (samara8954):

Last month, when Joe took his puppy to the veterinarian, the puppy weighed 4.7 pounds. This month, the puppy weighed 7.05 pounds. What is the percentage increase of the weight of the puppy?

OpenStudy (samara8954):

50% 33% 2% 67%

OpenStudy (anonymous):

33

OpenStudy (anonymous):

find what percentage is 7.05 from 4.7. 7.05/4.7*100%=150%

OpenStudy (samara8954):

is the answer 50

OpenStudy (samara8954):

yes or no

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@dtuniyants can you help me?

OpenStudy (anonymous):

since there is 50%, the answer is easily checked. 4.7*50%=2.35 4.7+2.35=7.05 ok

OpenStudy (anonymous):

what is your question @ms.hi

OpenStudy (samara8954):

The value of a professional basketball player's autograph rose 20% in the last year. It is now worth $276.00. What was it worth a year ago? $210.00 $220.00 $240.00 $230.00

OpenStudy (anonymous):

oel used candle molds, as shown, to make candles that were perfect cylinders and spheres What is the approximate difference in the amount of wax needed to make a candle from each of these molds? (Use π = 3.14.):

OpenStudy (anonymous):

4.19 10.4 12.56 20.93

OpenStudy (anonymous):

@ms.hi there are too lot of questions for me. I'll ask @AlexandervonHumboldt2 to answer you

OpenStudy (alexandervonhumboldt2):

i dont know

OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle if each of the two legs is 6 units?

OpenStudy (anonymous):

@Samara8954 use again this formula: new/old*100%=20%

OpenStudy (anonymous):

hypotenuse^2=(leg1)^2+(leg2)^2

OpenStudy (anonymous):

if two legs are 6 units. then hypotenuse is sqrt(6^2+6^2)=sqrt(36+36)=sqrt72 simplify this

OpenStudy (anonymous):

72 sqrt

OpenStudy (anonymous):

no i meant 12 sqrt

OpenStudy (anonymous):

sqrt72=sqrt(36*2)=6*sqrt2

OpenStudy (anonymous):

so its 6